A heating coil of resistance 20 ohms connected to a 200v source is used to boil a certain quantity of water in a container of heat capacity 100JK¹ for 2 minutes. of the initial temperature of the water is 40⁰c. Calculate th mass of the water on the container. (Specific heat capacity of water=4.2*10³Kg¹k¹, assume boiling point of watet=100⁰c)
1
Expert's answer
2020-08-12T16:45:06-0400

According to the Joule–Lenz law, the energy delivered to the system from the coil is:


"Q_{in} = \dfrac{U^2}{R}t"

where "U = 200V", "R = 20\space Ohms" and "t = 2min = 120s".

Thus, obtain:


"Q_{in} = \dfrac{200^2}{20}\cdot 120 = 2.4\times 10^{5}J"

The equation of the thermal balance will be:


"Q_{in} = C_{container}\Delta T + c_{water}m_{water}\Delta T"

where "C_{container} = 100J\/K" is the heat capacity of the container, "c_{water} = 4.2\times 10^3 J/kg\cdot K" is the specific heat capacity of water, "\Delta T = 100\degree C-40\degree C = 60\degree C = 60 K" is the temperature difference and "m_{water}" is the mass of water.

Expressing "m_{water}" from the last equation, obtain:


"m_{water} = \dfrac{Q_{in} - C_{container}\Delta T}{c_{water}\Delta T} = \dfrac{2.4\times 10^5 - 100\cdot 60}{4.2\times 10^3\cdot 60} \approx 0.929 kg"

Answer. 0.929 kg.